OpenStudy (anonymous):

find the limit x-->0 (x+1)/(x^2(x+7))

OpenStudy (xapproachesinfinity):

is that \(\large \lim_{x\to o}\frac{x+1}{x^2(x+7)}\)

OpenStudy (xapproachesinfinity):

@JonC

OpenStudy (anonymous):

@JonC: What kind of a question is this? I don't get it!

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

if so factor x^2 from top so you get x^2(1/x +1/x^2) then cancel x^2 and evaluate the limit after

OpenStudy (anonymous):

find the limit ?

OpenStudy (xapproachesinfinity):

Okay see what i said!

OpenStudy (xapproachesinfinity):

you need to factor x^2 from the top so you can cancel it

OpenStudy (xapproachesinfinity):

Do you get it?

OpenStudy (anonymous):

no sorry

OpenStudy (xapproachesinfinity):

Oh hold a second my bad I'm thinking about x===>OO for the top lol

OpenStudy (xapproachesinfinity):

Ok forgot what i said lol you need to do left and right limit

OpenStudy (xapproachesinfinity):

for the left and right limits you just plugging zero and think about the left side of zero or right side of zero

OpenStudy (xapproachesinfinity):

for the left one you should get 0+1/0(0+7) that x^2 will turn any small number near zero be positive so you get in the bottom 0+ so you got 1/0+ which is +OO

OpenStudy (anonymous):

@xapproachesinfinity you're talking crap stop it

OpenStudy (xapproachesinfinity):

then you do the same thing with right side limit 0+ the first one was 0- and changed to 0+ because we are squaring x^2

OpenStudy (xapproachesinfinity):

that's not how you should talk! if this is not relevant then say it in a nice way and offer a better way

OpenStudy (anonymous):

Or...... you're just talking sense.

OpenStudy (anonymous):

Fine you're talking "sense". There. Better now? You did a great job on the response though.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

sorry for acting mean to you @xapproachesinfinity you did good

OpenStudy (xapproachesinfinity):

here is the picture @JonC first limit from the left \(\large \lim_{x\to0^{-}}\frac{x+1}{x^2(x+7)}=\lim_{x\to0^-}\frac{0+1}{0^+(0+7)}\)

OpenStudy (xapproachesinfinity):

No worries ^^

OpenStudy (xapproachesinfinity):

so \(\large \lim_{x\to o-}\frac{1}{0^+}=\infty\) you do the same with right side limit

OpenStudy (xapproachesinfinity):

like i said when you evaluate small number close to zero and negative like for example -0.001 -0.001+7= gives a positive number and (-0.001)^2=positive as well so you get a number very close to zero but positive that's why we right it as 1/0+

OpenStudy (xapproachesinfinity):

Hope that makes sense to you! just something i want to throw here 1/0+=oo think about powers of then in the bottom like this \(10^{-2}\) so surely \(1/10^{-2}=10^2\) the more you get close to zero the more 1/0 blows to infinity